# Comparison of modularity-based approaches for nodes clustering in binary hypergraphs

### 25 Jan 2024 v1

@@ 3.2 Quality assessment
We first consider accuracy of the clustering, relying on the Adjusted Rand Index (ARI, Hubert and Arabie, 1985) that measures similarity between $\hat{\mathcal{C}}$ and $\mathcal{C}^{\textrm{true}}$ (up to label switching). It is upper bounded by 1, where a value of 1 indicates perfect agreement between the clusterings, and negative values indicate less agreement than expected by chance. Then we consider running times (expressed in seconds) of each method. The results have been obtained on a computer with a AMD EPYC 7542 32-Core processor, 128 CPU (2 sockets of 32 double threads cores; we used just one core for each job as none of the procedure is parallelized) and 675Gb RAM. We already mentioned that modularity maximization is far from trivial because of the size of the search space. Thus, an important question is whether the method at stake indeed maximizes its objective. To assess this, we measure the absolute relative error between the resulting value $\hat{Q}=Q(H,\hat{\mathcal{C}})$ at the estimated classification $\hat{\mathcal{C}}$ and the ground truth modularity $Q^{\textrm{true}}=Q(H,\mathcal{C}^{\textrm{true}})$, namely
$\text{error}=\frac{|Q(H,\hat{\mathcal{C}})-Q(H,\mathcal{C}^{\textrm{true}})|}{Q(H,\mathcal{C}^{\textrm{true}})}.$
A method that reaches its objective (modularity maximization) without being able to recover the true modular clusters would reveal that it is based on a definition of modularity that is not appropriate.
@@
Absolute relative error between $\hat{Q}$ and $Q^{\textrm{true}}$
@@ 4 Results
We first analyze the results under the simplest scenarios (namely scenarios A, which represent our base case) and the HSBM generating model. Results are presented in Figure 2. First, the CNM-like algorithm does not recover the ground truth clusters, with ARI values around 0 (Figure 2, top left). In fact, the algorithm did not improve over its initialization at $\mathcal{C}^{\text{own}}=(\{1\},\dots,\{n\})$ and the number of estimated clusters corresponds to the actual number of nodes (bottom right). Its absolute relative error on modularity is constant and corresponds to the absolute relative difference between the modularity of $\mathcal{C}^{\text{own}}$ and that of $\mathcal{C}^{\text{true}}$. The other 3 methods successfully recover the true clusters. For those 3 methods, median ARI values are above 0.7 (top left) and the number of estimated clusters varies between 3 and 6 (bottom right). While the AON-HMLL globally obtains the best ARI results (top left), it is also the fastest method (top right) and it attains its objective of modularity maximization (absolute relative error around 0, see bottom left). The LSR algorithm was proposed to improve over the IRMM and indeed its absolute relative error on modularity is (in general) improved over the latter (bottom left) with a much smaller dispersion. However from the clustering point of view, ARI is not improved (top left) and computing times are much larger (top right). This seems to indicate that the LSR places too much emphasis on maximizing modularity at the expense of clustering recovery. As the number of nodes $n$ increases, we observe that ARI values globally have a lower dispersion, but do not seem to overall improve (top left). This might be due to our setting where the within-cluster over between-cluster hyperedge ratio ($\rho_{s}$) is kept constant when $n$ varies.
Figure 2: Datasets HSBM, scenarios A1 to A5. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). The IRMM and (consequently) the LSR methods both gave an error on one dataset in scenario A5. Outlier points have been removed: from the absolute relative error plot (bottom left), 1 value above 500 concerning the IRMM method in scenario A1. Moreover, one dataset from scenario A5 gave an error with the IRMM and (consequently) the LSR methods; corresponding results were removed from the plots.
@@
Figure 3: Datasets DCHSBM, scenarios A1 to A6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario A6 range between 15,796 and 22,350 seconds and are not shown. Outlier points have been removed from the absolute relative error plot (bottom left): 1 value above 300 concerning the IRMM method in scenario A4.
To finish with these settings from scenarios A, we consider Figure 4 where the results for hypergraphs generated under the h-ABCD benchmark method are provided. Let us recall that while we tried to mimic as much as possible the characteristics of the scenarios A obtained under HSBM and DCHSBM, it was impossible to obtain similar node degrees within that h-ABCD generating process (see Table 3) and the ones obtained here are much smaller. We observe that in this setting, none of the proposed methods is able to reconstruct the true clusters: ARI values are generally lower than 0.3 (see Figure 4, top left) and the number of estimated clusters is too large (bottom right). Nonetheless, the modularity maximization seems to work as the absolute relative error between the ground truth modularity and its estimation is small (bottom left). This raises the following question: are these datasets indeed modular? We will come back to this later when discussing scenarios Z.
Figure 4: Datasets h-ABCD, scenarios A1 to A6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). Outlier points have been removed: from the absolute relative error plot (bottom left), 3 values at 25, 50 and 55 seconds concerning the IRMM method with in scenarios A6, A2 and A3 respectively.
@@
In scenario B, we decreased the sparsity wrt to scenario A (note that the hypergraphs remain nonetheless sparse, see Table 3). Results are presented in Figure 5. Here again, we removed from the time plot (top right) all values for the LSR method in scenario B6. Their range between 16,961 seconds and 17,895 seconds would have changed the $y$-scale. We mostly observe that while the above conclusions are still valid, the performances of the 3 “working” methods (AON-HMLL, IRMM and LSR) increase wrt to scenario A. Indeed, except for the CNM-like algorithm, the methods exactly recover the true number of clusters (bottom right) and ARI values are almost equal to 1 (top left). Absolute relative errors on modularity are also almost zero for those 3 methods, indicating that the local maximization of the modularity works. We note that the CNM-like method has absolute relative error equal to 1. This comes from the fact that the maximized modularity is zero while the ground truth modularity is not zero. Also note that the computing time for this method in scenario B6 becomes significantly larger.
Figure 5: Datasets DCHSBM, scenarios B1 to B6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario B6 range between 16,961 and 17,895 seconds and are not shown.
@@
Figure 6: Datasets HSBM, scenarios C1 to C5. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario C5 range between 2,646 and 3,842 seconds and are not shown. Outlier points have been removed: from the absolute relative error plot (bottom left), 2 values concerning the IRMM method, one above 30 in scenario C3 and the second above 60 in scenario C4.
@@
In scenario D, we explore the impact of the proportions of size-$s$ hyperedges. More precisely, while scenario A relied on a realistic setting of a smaller absolute number of size-$3$ than size-2 hyperedges (namely, $|\mathcal{E}_{2}|\gg|\mathcal{E}_{3}|$), we explore here the converse setting where $|\mathcal{E}_{3}|\gg|\mathcal{E}_{2}|$. Results are presented in Figure 7, where again, we removed from the time plot (top right), all values for the LSR method in scenario D6 for comparison purposes, as their range is between 17,153 and 22,377 seconds. Here, we observe as in scenario B that the performances of the 3 methods AON-HMLL, IRMM and LSR increase wrt to scenario A. Indeed, the AON-HMLL and the IRMM have ARI values equal or close to 1 (top left) and find (almost always) the correct number of clusters (bottom right), indicating (almost) perfect clustering recovery. Absolute relative errors on modularity are also almost zero for those 3 methods, indicating that the local maximization of the modularity works. From that simulation we conclude that clustering via modularity maximization is easier for datasets with a larger proportion of large-size hyperedges and conversely, more difficult in the realistic setting where larger sizes hyperedges are in smaller proportion.
Figure 7: Datasets DCHSBM, scenarios D1 to D6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario D6 range between 17,153 and 22,377 seconds and are not shown. Outlier points have been removed: from the time plot (top right), 1 value above 2,800 seconds concerning the LSR method in scenario D5 and from the absolute relative error plot (bottom left): 2 values larger than 10 concerning the IRMM method in scenario D5.
@@
Scenario E (resp. F) rely on a larger (resp. smaller) value for the within-cluster over between-cluster hyperedge ratio $\rho_{s}$ (still constant with hyperedge size $s$) compared to scenario A. The results of this simulation are presented in Figure 8) (resp. Figure 9). In those figures again, time values for the LSR method in scenarios E6 and F6 respectively have been removed. We can observe that the modularity based methods are sensitive to this parameter $\rho_{s}$, with better clustering results obtained when this ratio is large. As expected, the more modular the hypergraphs are, the easier it is to recover the clusters.
Figure 8: Datasets DCHSBM, scenarios E1 to E6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for LSR method in scenario E6 range between 15,833 and 20604 seconds and are not shown. Outlier points have been removed from the absolute relative error plot (bottom left): 2 values concerning the IRMM method, one larger than 100 in scenario E3 and the other larger than 20 in scenario E4.
Figure 9: Datasets DCHSBM, scenarios F1 to F6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for LSR method in scenario F6 range between 22,891 and 39,967 seconds and are not shown. Outlier points have been removed from the absolute relative error plot (bottom left): 3 values concerning the IRMM method, with values larger than 40, 680 and 20 in scenarios F1, F4 and F6 respectively.
@@
Figure 10: Datasets h-ABCD, scenarios Z1 to Z5. Comparison by increasing the number of nodes from $n\in\{100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), absolute relative error on modularity (bottom left) and difference between estimated number of clusters and true value (bottom right). Outlier points have been removed: from the time plot (top right), 4 values above 900 seconds concerning the LSR method in scenario Z5 and from the absolute relative error plot (bottom left), 1 value above 25 concerning the IRMM method in scenario Z4.
@@
As a final note, we notice that in our experiments, the IRMM method sometimes shows some very large values for the absolute relative modularity error (points that we called “outliers” and removed to preserve $y$-scales in the plots). This indicates a large difference in the values of the $Q^{\textrm{w-clique}}$ modularity at the ground truth clustering and at the maximized value.
@@ 5 Discussion
  • Though the LSR algorithm is specifically designed to improve on the IRMM, it does not improve the clustering problem at stake.
@@
  • We did not observe any algorithm for which the modularity $Q$ would be correctly maximized (relative absolute error in modularity close to zero) while clusters would not be recovered (low ARI values). Nonetheless, the modularity $Q^{\textrm{strict}}$ from Kamiński et al. (2019a) is not fully maximized by the CNM-like method, which leaves open the question of whether it is able to capture communities in hypergraphs.
@@
  • The methods tend to have better results when we observe a larger proportion of larger-size hyperedges (i.e., when $|\mathcal{E}_{3}|$ becomes larger than $|\mathcal{E}_{2}|$, see scenarios D).
@@
Another conclusion from our study is that the h-ABCD benchmark model (Kamiński et al., 2023b) does not seem appropriate to generate modular graphs, or at least that none of the current modularity-based approaches is able to detect the simulated clusters in those hypergraphs.
@@
Our work is a first building block in gaining a better understanding of modularity in hypergraphs, yet it comes with certain limitations that warrant attention in future research. One constraint arises from computational limitations in both the generating models and modularity maximization methods, restricting our exploration to relatively small graphs (with a number of nodes $n\leq 1,000$). Consequently, we constrained ourselves to a limited number of clusters ($K\leq 3$), as larger values might lead to clusters too small for effective detection. Our focus was on binary hypergraphs, which already encompass a vast array of higher-order interactions. However, weighted hypergraphs are also of significant interest. Additionally, our approach relied on simulated hypergraphs with characteristics dictated by methodological constraints (e.g., the number of nodes, number of clusters) and others chosen to align with what we believe to be realistic (e.g., sparse hypergraphs, $|\mathcal{E}_{2}|\gg|\mathcal{E}_{3}|$, $\dots$). In a related study Lee et al. (2021), the authors examined 13 real-world hypergraphs with heterogeneous sparsity (ratios $|\mathcal{E}|/n$ ranging from as small as 0.5 to around 50) and an average hyperedge size $s$ generally less than 3.9, with two exceptions (hypergraphs related to drug chemicals). Despite this attempt, the literature still lacks a large-scale study on the characteristics of real-world hypergraphs that could inform and support simulations.
@@
There remain numerous unresolved questions that extend beyond the scope of the present contribution. Realistic characteristics, influenced by parameter choices in the generating models, are intricately tied to the issue of detectability thresholds. Specifically, under what circumstances is it possible to effectively recover clusters in a hypergraph? While this question has garnered attention for uniform hypergraphs (Angelini et al., 2015; Chien et al., 2019; Stephan and Zhu, 2022; Zhang and Tan, 2023), real-world hypergraphs, which are non-uniform, remain largely unexplored in this context. Furthermore, moving beyond clustering recovery, it would be valuable to investigate the discriminative power of modularities. Specifically, understanding how discriminative each proposed modularity measure is could provide insights. Examining the distribution of modularity values across a diverse set of hypergraphs, including non-modular ones, holds significant importance. In a similar vein, our observation of cases where the IRMM method exhibited a very large absolute relative error on modularity raises questions about whether the corresponding modularity (and others) is unimodal. Characterizing the behavior of modularities across the entire spectrum of node clusterings would aid in designing suitable modularity-based methods for community detection in hypergraphs.

### 04 Mar 2024 v2

@@ 3.2 Quality assessment
We first consider accuracy of the clustering, relying on the Adjusted Rand Index (ARI, Hubert and Arabie, 1985) that measures similarity between $\hat{\mathcal{C}}$ and $\mathcal{C}^{\textrm{true}}$ (up to label switching). It is upper bounded by 1, where a value of 1 indicates perfect agreement between the clusterings, and negative values indicate less agreement than expected by chance. Then we consider running times (expressed in seconds) of each method. The results have been obtained on a computer with a AMD EPYC 7542 32-Core processor, 128 CPU (2 sockets of 32 double threads cores; we used just one core for each job as none of the procedure is parallelized) and 675Gb RAM. We already mentioned that modularity maximization is far from trivial because of the size of the search space. Thus, an important question is whether the method at stake indeed maximizes its objective. To assess this, we measure the relative error between the ground truth modularity $Q^{\textrm{true}}=Q(H,\mathcal{C}^{\textrm{true}})$ and the resulting value $\hat{Q}=Q(H,\hat{\mathcal{C}})$ at the estimated classification $\hat{\mathcal{C}}$, namely
$\text{error}=\frac{Q(H,\mathcal{C}^{\textrm{true}})-Q(H,\hat{\mathcal{C}})}{Q(H,\mathcal{C}^{\textrm{true}})}.$
A method that reaches its objective (modularity maximization) without being able to recover the true modular clusters would reveal that it is based on a definition of modularity that is not appropriate. Also note that this error has a sign, with negative values indicating that ground truth modularity is not the maximum value. The mean values and standard deviations for ground truth modularity $Q^{\textrm{true}}$ are also reported (Table 5 in Appendix), since values close to zero could induce unstable errors.
@@
Relative error between $Q^{\textrm{true}}$ and $\hat{Q}$
@@ 4 Results
We first analyze the results under the simplest scenarios (namely scenarios A, which represent our base case) and the HSBM generating model. Results are presented in Figure 2. First, the CNM-like algorithm does not recover the ground truth clusters, with ARI values around 0 (Figure 2, top left). In fact, the algorithm did not improve over its initialization at $\mathcal{C}^{\text{own}}=(\{1\},\dots,\{n\})$ and the number of estimated clusters corresponds to the actual number of nodes (bottom right). Its relative error on modularity is constant and corresponds to the relative difference between the modularity of $\mathcal{C}^{\text{true}}$ and that of $\mathcal{C}^{\text{own}}$. It is positive, so that the modularity maximization goal is clearly not achieved here. The other 3 methods successfully recover the true clusters. For those 3 methods, median ARI values are above 0.7 (top left) and the number of estimated clusters varies between 3 and 6 (bottom right). While the AON-HMLL globally obtains the best ARI results (top left), it is also the fastest method (top right) and it attains its objective of modularity maximization (relative error around 0, see bottom left). The LSR algorithm was proposed to improve over the IRMM. While its relative error on modularity (bottom left) seems in general improved over the latter (with smaller values), Table 5 in Appendix shows that the modularity $Q^{\textrm{w-clique}}$ optimized by IRMM is close to 0 for the ground truth clusters, thus giving unstable errors; while $Q^{\textrm{linear}}$ optimized by LSR is strictly positive at those ground truth clusters. Most importantly, from the clustering point of view, ARI is not improved (top left) and computing times are much larger (top right). This seems to indicate that the LSR places too much emphasis on maximizing modularity at the expense of clustering recovery. As the number of nodes $n$ increases, we observe that ARI values globally have a lower dispersion, but do not seem to overall improve (top left). This might be due to our setting where the within-cluster over between-cluster hyperedge ratio $\rho_{s}$ is kept constant when $n$ varies.
Figure 2: Datasets HSBM, scenarios A1 to A5. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). The IRMM and (consequently) the LSR methods both gave an error on one dataset in scenario A5. Outlier points have been removed: from the relative error plot (bottom left), 1 value below -500 concerning the IRMM method in scenario A1. Moreover, one dataset from scenario A5 gave an error with the IRMM and (consequently) the LSR methods; corresponding results were removed from the plots.
@@
Figure 3: Datasets DCHSBM, scenarios A1 to A6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario A6 range between 15,796 and 22,350 seconds and are not shown. Outlier points have been removed from the relative error plot (bottom left): 1 value above 300 concerning the IRMM method in scenario A4.
To finish with these settings from scenarios A, we consider Figure 4 where the results for hypergraphs generated under the h-ABCD benchmark method are provided. Let us recall that while we tried to mimic as much as possible the characteristics of the scenarios A obtained under HSBM and DCHSBM, it was impossible to obtain similar node degrees within that h-ABCD generating process (see Table 3) and the ones obtained here are much smaller. We observe that in this setting, none of the proposed methods is able to reconstruct the true clusters: ARI values are generally lower than 0.3 (see Figure 4, top left) and the number of estimated clusters is too large (bottom right). Nonetheless, the modularity maximization seems to work as the relative error between the ground truth modularity and its estimation is small (bottom left). Note also that the LSR algorithm seems to find a clustering with larger value of the $Q^{\textrm{linear}}$ modularity than at the ground truth clusters (negative errors). Overall, our conclusions raise the following question: are these datasets indeed modular? We will come back to this later when discussing scenarios Z.
Figure 4: Datasets h-ABCD, scenarios A1 to A6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). Outlier points have been removed: from the relative error plot (bottom left), 3 values at 25, -50 and -55 concerning the IRMM method with in scenarios A6, A2 and A3 respectively.
@@
In scenario B, we decreased the sparsity wrt to scenario A (note that the hypergraphs remain nonetheless sparse, see Table 3). Results are presented in Figure 5. Here again, we removed from the time plot (top right) all values for the LSR method in scenario B6. Their range between 16,961 seconds and 17,895 seconds would have changed the $y$-scale. We mostly observe that while the above conclusions are still valid, the performances of the 3 “working” methods (AON-HMLL, IRMM and LSR) increase wrt to scenario A. Indeed, except for the CNM-like algorithm, the methods exactly recover the true number of clusters (bottom right) and ARI values are almost equal to 1 (top left). Relative errors on modularity are also almost zero for those 3 methods, indicating that the local maximization of the modularity works. We note that the CNM-like method has relative error equal to 1. This comes from the fact that the maximized modularity is zero while the ground truth modularity is not zero. Also note that the computing time for this method in scenario B6 becomes significantly larger.
Figure 5: Datasets DCHSBM, scenarios B1 to B6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario B6 range between 16,961 and 17,895 seconds are not shown.
@@
Figure 6: Datasets HSBM, scenarios C1 to C5. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario C5 range between 2,646 and 3,842 seconds and are not shown. Outlier points have been removed: from the relative error plot (bottom left), 2 values concerning the IRMM method, one above 30 in scenario C3 and the second below -60 in scenario C4.
@@
In scenario D, we explore the impact of the proportions of size-$s$ hyperedges. More precisely, while scenario A relied on a realistic setting of a smaller number of size-$3$ than size-2 hyperedges (namely, $|\mathcal{E}_{2}|\gg|\mathcal{E}_{3}|$), we explore here the converse setting where $|\mathcal{E}_{3}|\gg|\mathcal{E}_{2}|$. Results are presented in Figure 7, where again, we removed from the time plot (top right), all values for the LSR method in scenario D6 for comparison purposes, as their range is between 17,153 and 22,377 seconds. Here, we observe as in scenario B that the performances of the 3 methods AON-HMLL, IRMM and LSR increase wrt to scenario A. Indeed, the AON-HMLL and the IRMM have ARI values equal or close to 1 (top left) and find (almost always) the correct number of clusters (bottom right), indicating (almost) perfect clustering recovery. Relative errors on modularity are also almost zero for those 3 methods, indicating that the local maximization of the modularity works. From that simulation we conclude that clustering via modularity maximization is easier for datasets with a larger proportion of large-size hyperedges and conversely, more difficult in the realistic setting where larger sizes hyperedges are in smaller proportion.
Figure 7: Datasets DCHSBM, scenarios D1 to D6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for the LSR method in scenario D6 range between 17,153 and 22,377 seconds and are not shown. Outlier points have been removed: from the time plot (top right), 1 value above 2,800 seconds concerning the LSR method in scenario D5 and from the relative error plot (bottom left): 1 value larger than 14 and 1 smaller than -11 concerning the IRMM method in scenario D5.
@@
Scenario E (resp. F) rely on a larger (resp. smaller) value for the within-cluster over between-cluster hyperedge ratio $\rho_{s}$ (still constant with hyperedge size $s$) compared to scenario A. The results of this simulation are presented in Figure 8 (resp. Figure 9). In those figures again, time values for the LSR method in scenarios E6 and F6 respectively have been removed. We can observe that the modularity based methods are sensitive to this parameter $\rho_{s}$, with better clustering results obtained when this ratio is large. As expected, the more modular the hypergraphs are, the easier it is to recover the clusters.
Figure 8: Datasets DCHSBM, scenarios E1 to E6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for LSR method in scenario E6 range between 15,833 and 20604 seconds and are not shown. Outlier points have been removed from the relative error plot (bottom left): 2 values concerning the IRMM method, one larger than 100 in scenario E3 and the other smaller than -21 in scenario E4.
Figure 9: Datasets DCHSBM, scenarios F1 to F6. Comparison by increasing the number of nodes from $n\in\{50,100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and estimated number of clusters (bottom right, true value is $3$). From the time plot (top right), values for LSR method in scenario F6 range between 22,891 and 39,967 seconds and are not shown. Outlier points have been removed from the relative error plot (bottom left): 3 values concerning the IRMM method, with 2 values smaller than -40 and -680 and 1 value larger than 22 in scenarios F1, F4 and F6 respectively.
@@
Figure 10: Datasets h-ABCD, scenarios Z1 to Z5. Comparison by increasing the number of nodes from $n\in\{100,150,200,500,1000\}$: Adjusted Rand Index (top left), time in seconds (top right), relative error on modularity (bottom left) and difference between estimated number of clusters and true value (bottom right). Outlier points have been removed: from the time plot (top right), 4 values above 900 seconds concerning the LSR method in scenario Z5 and from the relative error plot (bottom left), 1 value below -28 concerning the IRMM method in scenario Z4.
@@
As a final note, we notice that in our experiments, the IRMM method sometimes shows some very large values for the relative modularity error (points that we called “outliers” and removed to preserve $y$-scales in the plots). Looking at Table 5 in the Appendix we observe that the ground truth modularity $Q^{\textrm{w-clique}}$ is close to zero, explaining this unstable behaviour of the relative error.
@@ 5 Discussion
  • Though the LSR algorithm is specifically designed to improve on the IRMM, it does not improve the clustering problem at stake;
@@
  • We did not observe any algorithm for which the modularity $Q$ would be correctly maximized (relative error in modularity close to zero) while clusters would not be recovered (low ARI values). Nonetheless, the modularity $Q^{\textrm{strict}}$ from Kamiński et al. (2019a) is not fully maximized by the CNM-like method, which leaves open the question of whether it is able to capture communities in hypergraphs.
@@
  • The methods tend to have better results when we observe a larger proportion of larger-size hyperedges (i.e., when $|\mathcal{E}_{3}|$ becomes larger than $|\mathcal{E}_{2}|$, see scenarios D);
@@
Another conclusion from our study is that the h-ABCD benchmark model (Kamiński et al., 2023b) does not seem appropriate to generate modular hypergraphs, or at least that none of the current modularity-based approaches is able to detect the simulated clusters in those hypergraphs.
@@
Our work is a first building block in gaining a better understanding of modularity in hypergraphs, yet it comes with certain limitations that warrant attention in future research. One constraint arises from computational limitations in both the generating models and modularity maximization methods, restricting our exploration to relatively small graphs (with a number of nodes $n\leq 1,000$). Consequently, we constrained ourselves to a limited number of clusters ($K\leq 3$), as larger values might lead to clusters too small for effective detection. Our focus was on binary hypergraphs, which already encompass a vast array of higher-order interactions. However, weighted hypergraphs are also of significant interest. Additionally, our approach relied on simulated hypergraphs with characteristics dictated by methodological constraints (e.g., the number of nodes, number of clusters) and others chosen to align with what we believe to be realistic (e.g., sparse hypergraphs, $|\mathcal{E}_{2}|\gg|\mathcal{E}_{3}|$, $\dots$). Lee et al. (2021) examined 13 real-world hypergraphs with heterogeneous sparsity (ratios $|\mathcal{E}|/n$ ranging from as small as 0.5 to around 50) and an average hyperedge size $s$ generally less than 3.9, with two exceptions (hypergraphs related to drug chemicals). Despite this attempt, the literature still lacks a large-scale study on the characteristics of real-world hypergraphs that could inform and support simulations.
@@
There remain numerous unresolved questions that extend beyond the scope of the present contribution. Realistic characteristics, influenced by parameter choices in the generating models, are intricately tied to the issue of detectability thresholds. Specifically, under what circumstances is it possible to effectively recover clusters in a hypergraph? While this question has garnered attention for uniform hypergraphs (Angelini et al., 2015; Chien et al., 2019; Stephan and Zhu, 2022; Zhang and Tan, 2023), real-world hypergraphs, which are non-uniform, remain largely unexplored in this context. Furthermore, moving beyond clustering recovery, it would be valuable to investigate the discriminative power of modularities. Specifically, understanding how discriminative each proposed modularity measure is could provide insights on their design. Examining the distribution of modularity values across a diverse set of hypergraphs, including non-modular ones, holds significant importance. In a similar vein, whether hypergraph modularities are unimodal or not is an important question. Characterizing the behavior of modularities across the entire spectrum of node clusterings would aid in designing suitable modularity-based methods for community detection in hypergraphs.